OpenStudy (anonymous):

You sell sporting goods. Your wages depend on the value of your sales. One week you sold $3,500 in sporting goods, earning, $950. Another week you sold $2800 in sporting goods, earning $810. Write an equation of the line in standard form

OpenStudy (anonymous):

Slope= (3500-2800) / (950-810) =700/ 140 =5 y=mx+b is the formula for a line y=5x+b substitute 3500=5*950+b b=-1250 y=5x-1250